# FairTranslate: An English-French Dataset for Gender Bias Evaluation in Machine Translation by Overcoming Gender Binarity

### v1

@@ Abstract.
ccs: Computing methodologies Machine translation
@@ 3. Related Work
Evaluating gender bias in Machine Translation (MT) has been extensively studied using benchmarks like Winogender (Rudinger et al., 2018), Winobias (Zhao et al., 2018), and WinoMT (Stanovsky et al., 2019), which assess coreference resolution to determine how models attribute gender based on context. Similarly, MT-GenEval (Currey et al., 2022) employs a counterfactual methodology by systematically modifying sentences to analyze gender biases. Our approach integrates both coreference resolution and counterfactual evaluation, extending these methodologies beyond the binary framework to include non-binary gender and address the broader challenges associated with it.
Recent works addressing non-binary gender, such as WinoQueer Felkner et al. (2023) and Dhingra et al. (2023), focus on evaluating bias related to gender and LGBTQ+ identities but do not specifically address machine translation. AmbGIMT (Chen et al., 2024) evaluates non-binary gender in English-Chinese MT but targets attitude translation rather than coreference or singular ’they’. In spite of this, having both English-Chinese and English-French datasets opens up new opportunities in exploring machine translation of non-binary gender terms across vastly different languages such as French and Chinese.

### v2

@@ Abstract.
journalyear: 2025copyright: acmlicensedconference: The 2025 ACM Conference on Fairness, Accountability, and Transparency; June 23–26, 2025; Athens, Greecebooktitle: The 2025 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’25), June 23–26, 2025, Athens, Greecedoi: 10.1145/3715275.3732013isbn: 979-8-4007-1482-5/2025/06ccs: Computing methodologies Machine translation
@@ 3. Related Work
Evaluating gender bias in Machine Translation (MT) has been extensively studied using benchmarks like Winogender (Rudinger et al., 2018), Winobias (Zhao et al., 2018), WinoMT (Stanovsky et al., 2019) and, more recently, WinoPron (Gautam et al., 2024), which assess coreference resolution to determine how models attribute gender based on context. Similarly, MT-GenEval (Currey et al., 2022) employs a counterfactual methodology by systematically modifying sentences to analyze gender biases. Our approach integrates both coreference resolution and counterfactual evaluation, extending these methodologies beyond the binary framework to include non-binary gender and address the broader challenges associated with it.
Recent studies addressing non-binary gender (Felkner et al., 2023; Dhingra et al., 2023; Waldis et al., 2024) focus on evaluating bias related to gender and LGBTQ+ identities, but do not specifically target machine translation. AmbGIMT (Chen et al., 2024) evaluates non-binary gender in English-Chinese MT, but its focus is on attitude translation rather than coreference or the use of singular they. An English-German MT study also exists, focusing on gender-neutral person-referring terms (Lardelli et al., 2024).
While these works cover English-Chinese and English-German, the addition of an English-French dataset enables exploration of non-binary gender translation across typologically diverse languages such as French, German, and Chinese, opening up new research opportunities.